OpenStudy (anonymous):

ah! confusing one 1.12=1.25g=8.62

jhonyy9 (jhonyy9):

this is confused you realy

OpenStudy (anonymous):

yeah and i typed it wrong that was supposed to be a + by 1.12

OpenStudy (anonymous):

try writing it in fractions, i think that shud help

OpenStudy (anonymous):

1.12=28/25 1.25=5/4

OpenStudy (anonymous):

wait a sec u said u typed it in wrong?

OpenStudy (anonymous):

yes i meant a + not an =

OpenStudy (anonymous):

i got 6?

OpenStudy (anonymous):

oh ok screw the fractions then lol my bad subtract 1.12 from both sides and u get 1.25g= 7.5 then divide 1.25 on both sides and u get 6

OpenStudy (anonymous):

so u r right!!

OpenStudy (anonymous):

lol! thank you!

OpenStudy (anonymous):

np :)